OpenStudy (anonymous):

How many times more light does a 4 meter mirror collect than a 1 meter mirror? Assume both mirrors are circular

OpenStudy (kropot72):

The area of the mirror determines how much light it will collect. The area of a circle is given by \[A _{circle}=\frac{\pi d ^{2}}{4}\] So the area is proportional to the diameter squared. So the increase in light that is collected will be \[\frac{4^{2}}{1^{2}}=?\]

OpenStudy (anonymous):

So would it collect 16 times more light than the 1 meter?

OpenStudy (kropot72):

Yes, you are correct.

OpenStudy (anonymous):

Thank you so much!

OpenStudy (kropot72):

You're welcome :)